# An Efficient Row-Based Sparse Fine-Tuning

### v1

@@ Abstract
Fine-tuning is an important step in adapting foundation models such as large language models to downstream tasks. To make this step more accessible to users with limited computational budgets, it is crucial to develop fine-tuning methods that are memory and computationally efficient. Sparse Fine-tuning (SFT) and Low-rank adaptation (LoRA) are two frameworks that have emerged for addressing this problem and have been adopted widely in practice. In this work, we develop a new SFT framework, based on ideas from neural network pruning. At a high level, we first identify ”important” neurons/nodes using feature importance metrics from network pruning (specifically, we use the structural pruning method), and then perform fine-tuning by restricting to weights involving these neurons. Using experiments on common language tasks, we demonstrate that our method significantly improves the memory efficiency of SFT without increasing training time complexity and implementation complexity, while achieving accuracy comparable to state-of-the-art methods such as LoRA and its variants.
Efficient AI, transfer learning, foundation model.
@@ 1 Introduction
Despite the success, fine-tuning a model with billions of parameters requires access to heavy computational resources, even when the task datasets are fairly small. Fortunately, studies (e.g., (Panigrahi et al., 2023) and references therein) show that fine-tuning only a small fraction of parameters can be effective. Parameter-efficient fine-tuning (PEFT) methods have thus been proposed to carry out this idea and address the challenge of making fine-tuning more accessible (Lialin et al., 2023). A leading PEFT approach, Low-Rank Adaptation (LoRA, Hu et al. 2022), achieves memory efficiency by simply making low-rank updates to the weight matrices in the different layers. Another class of PEFT methods is sparse fine-tuning (SFT, Sung et al. 2021; Guo et al. 2021; Ansell et al. 2022; Nikdan et al. 2024), which learns a sparse matrix, typically an unstructured one, for updating the pre-trained weights. However, SFT typically incurs higher memory costs than LoRA during the fine-tuning process, because of the unstructured sparsity. Several works aim to mitigate the memory complexity of SFT (Mofrad et al., 2019; Holmes et al., 2021; Nikdan et al., 2023, 2024), often at the cost of increased running time and more complex implementations of sparse kernels. Besides PEFTs, techniques like Zeroth-Order optimization (Malladi et al., 2023; Guo et al., 2024b) and quantization (Gholami et al., 2022; Dettmers et al., 2022, 2024) can further enhance memory and training efficiency for fine-tuning, including LoRA and SFT.
As LLMs increase in scale, advancing efficient sparse matrix computation, PEFT, and efficient training remains a crucial problem. Towards this goal, we study the question: Can sparse fine-tuning be improved to create a memory- and parameter-efficient framework, while avoiding additional implementations of sparse operations and without increasing the training time complexity? We answer this question in the affirmative, by proposing a new SFT framework for fine-tuning LLMs and Vision Transformers that achieves memory- and parameter-efficiency while maintaining or even improving performance on downstream tasks. Our approach utilizes NN pruning techniques to identify a subset of fine-tuning parameters and employs a matrix decomposition-based computation for efficient fine-tuning. This design enables the integration of ideas from model compression, SFT, and matrix decomposition methods.
@@ 1.1 Our Contributions
  • We leverage ideas from network pruning to improve SFT, achieving significant memory efficiency considerably lower than the popular LoRA. Our method uses only standard tensor operations, eliminating the need for custom sparse tensor libraries. Our approach is also modular, and it allows us to integrate several existing pruning techniques (which give different neuron importance scores) and works with all layer types, including LayerNorm and BatchNorm, which LoRA cannot directly handle.
  • We analyze the memory assignment of several PEFT methods and suggest that model architecture and computation graphs affect memory more significantly than the number of trainable parameters. We validate our methods across diverse fine-tuning tasks (language and vision) and provide practical guidance on training strategies to maximize efficiency and accuracy.
  • We propose two variants of the Taylor importance for different settings in image and language tasks: class-aware Taylor and Zeroth-Order Taylor. The first one is designed for tasks where class-wise accuracy is important (in addition to overall accuracy), such as image classification. Zeroth-Order Taylor is designed for large language models and requires memory only equal to that of a forward pass. In addition, we show how to effectively reduce the estimation variance of the Zeroth-Order estimator.
@@ 2 Background and Related Work
Parameter-Efficient and Memory-Efficient Fine-Tuning: In various language and vision tasks, the “pre-train then fine-tune” paradigm has been shown highly effective. PEFT methods (Lialin et al., 2023) fine-tune a small subset of the parameters of a large pre-trained model in order to accelerate the training process. We begin by introducing SFT and LoRA, two popular approaches for PEFT.
Sparse Fine-Tuning: SFT formulates the fine-tuning process as learning another weight matrix ${\mathbf{W}}_{s}$:
$\hat{{\mathbf{W}}}={\mathbf{W}}+{\mathbf{W}}_{s},$ (1)
${\mathbf{h}}=f(\hat{{\mathbf{W}}},{\mathbf{x}})=f({\mathbf{W}}+{\mathbf{W}}_{s},{\mathbf{x}}),$ (2)
where ${\mathbf{h}}\in\mathbb{R}^{d_{out}}$ and ${\mathbf{x}}\in\mathbb{R}^{d_{in}}$ are the input and output of the hidden layer, respectively, $f(\cdot)$ is the forward function, ${\mathbf{W}}\in\mathbb{R}^{d_{out}\times d_{in}}$ represents the frozen pre-trained parameters, and $\hat{{\mathbf{W}}}\in\mathbb{R}^{d_{out}\times d_{in}}$ denotes the final parameters used during inference for the fine-tuning task. The matrix ${\mathbf{W}}_{s}\in\mathbb{R}^{d_{out}\times d_{in}}$ is sparse and is initialized at ${\bm{0}}$. Such a decomposition is done for every layer in the neural network. SFT methods try to limit the number of parameters to fine-tune. For selecting non-zero indices, Diff pruning (Guo et al., 2021) learns a mask for ${\mathbf{W}}_{s}$ (using a standard Backprop algorithm), while FISH Mask (Sung et al., 2021) uses Fisher information to identify important indices in ${\mathbf{W}}$. Lottery Ticket SFT (Ansell et al., 2022) fine-tune the whole model for one epoch, then use ${\mathbf{W}}_{s}$ itself as an importance score to decide which parameters to fine-tune subsequently. Robust Adaptor (RoSA, Nikdan et al. 2024) combines the above SFTs with LoRA and outperforms all these approaches. However, the key challenge of all SFT methods is that they do not sufficiently reduce memory usage, as ${\mathbf{W}}_{s}$ keeps the dimensionality of ${\mathbf{W}}$, and thus standard libraries do not yield memory improvements.
@@
However, these techniques come with trade-offs, particularly in terms of time complexity and implementation complexity. Achieving memory efficiency often requires a significant increase in time complexity. To mitigate this, some approaches employ optimizations implemented in C++ or lower-level languages, such as those used in (Gale et al., 2020; Nikdan et al., 2023, 2024), to accelerate the training process.
Low-Rank Adaptation (LoRA): Instead of requiring ${\mathbf{W}}_{s}$ to be sparse, low-rank adaptation aims to find update matrices that are of small rank:
@@
where $\alpha$ is the LoRA scaling hyper-parameter, ${\mathbf{B}}\in\mathbb{R}^{d_{out}\times r},\ {\mathbf{A}}\in\mathbb{R}^{r\times d_{in}}$ are the low-rank matrices with $r\ll d_{in},d_{out}$. During inference, the ${\mathbf{B}}{\mathbf{A}}$ term can be merged into ${\mathbf{W}}$ to maintain the inference latency of the original model. During training, owing to the fact that $f$ is additive for both the self-attention blocks and the subsequent multilayer perceptron (MLP) layers of transformers (Vaswani, 2017), backpropagation can be performed efficiently for the ${\mathbf{B}},{\mathbf{A}}$ parameters. Due to LoRA’s simplicity and effectiveness, numerous variants have been proposed to enhance the performance, e.g., QLoRA (Dettmers et al., 2022; Guo et al., 2024a; Li et al., 2024; Dettmers et al., 2024), DoRA (Liu et al., 2024), RoSA (Nikdan et al., 2024), and VeRA (Kopiczko et al., 2024). These methods have achieved exceptional performance, often comparable to full fine-tuning across a range of tasks.
@@ 3 Number of Trainable Parameters Is Not Everything
Before introducing our approach, we want to emphasize that in PEFT research, reducing the number of trainable parameters is not the most critical factor for saving memory. Some PEFT methods mentioned in Section 2 focus on reducing the number of trainable parameters to decrease memory consumption. However, once the number of trainable parameters is sufficiently small, this reduction is no longer the most critical factor influencing memory usage.
Llama2(7B)
PEFT #param mem dropout param other
LoRA 159.9M(2.37%) 23.46GB 6.18GB 2.51GB 14.77GB
RoSA 157.7M(2.34%) 44.69GB 6.22GB 1.72GB 36.75GB
RoSA-bf16 157.7M(2.34%) 39.55GB 6.22GB 0.92GB 32.41GB
DoRA 161.3M(2.39%) 44.85GB 6.29GB 2.49GB 36.07GB
VeRA 1.37M(0.02%) 22.97GB 2.15GB 0.11GB 20.71GB
Table 1: The requirements of computation resources for fine-tuning full precision Llama2 using LoRA (Hu et al., 2022), RoSA (Nikdan et al., 2024), DoRA (Liu et al., 2024), and VeRA (Kopiczko et al., 2024). In this analysis, we set $r=64$ for LoRA, DoRA, and VeRA and set $r=32,d=1.2\%$ for RoSA. The fine-tuning parameters of RoSA-bf16 are in bfloat16 under mixed-precision training (Micikevicius et al., 2018), while all others are in full precision. Note that RoSA has its own official implementation, whereas LoRA, DoRA, and VeRA are integrated into the PEFT library provided by Hugging Face. This may influence the memory used by each method. ‘mem’ traces the peak memory usage during training, excluding the memory consumed by the model itself, details please see Appendix C.3. ‘dropout’ accounts for the memory consumption associated with LoRA’s dropout layer. ‘param’ represents the maximum memory savings from reducing the number of trainable parameters. It is quantified by varying $r$ and $d$ for each method and calculating the memory difference attributed to trainable parameters. ‘other’ indicates the total memory usage from any other intermediate values in backpropagation and optimizer states.
In NN training, backpropagation involves caching numerous intermediate values to compute gradients efficiently for each tensor. The memory cost of these intermediate values is heavily influenced by the computation graph and the model architecture. When the number of trainable parameters is relatively small, the memory consumed by intermediate values far exceeds that of the trainable parameters themselves. We use VeRA, RoSA, and DoRA to demonstrate the influences of these factors. Table 1 provides a detailed breakdown of the memory usage across different PEFT methods and components. For ‘dropout’, we traced the memory usage of fine-tuning Llama2 using these PEFT methods, with and without dropout, to yield dropout’s memory cost. Additionally, we examined memory usage under varying $r$ and $d$ to obtain the maximum potential memory savings from reducing the number of trainable parameters in each method. For instance, in the case of LoRA, fine-tuning the model with $r=64$ and $r=32$ shows a memory saving of approximately $1.25$GB when reducing $r$ from $64$ to $32$. This implies that the memory consumed by trainable parameters in LoRA with $r=64$ is approximately $1.25\times 2=2.51$GB.
Among PEFT methods that build on LoRA, VeRA aims to further reduce memory by lowering the number of trainable parameters, while RoSA and DoRA focus on improving accuracy. VeRA shares a unique pair of low-rank matrices across layers, reducing trainable parameters compared to LoRA. However, as shown in Table 1, it saves about 2.4GB from trainable parameters and about 4GB from dropout layers while introducing extra overhead from shared parameters and the corresponding intermediate values, resulting in minimal overall savings. DoRA and RoSA consume significantly more memory due to their complex computation and unstructured sparse matrices. DoRA decomposes LoRA’s matrices into magnitude and direction (see figure 3 in Appendix C.4), which significantly increases memory usage. As reflected in Table 1, while the memory cost by trainable parameters in DoRA is similar to LoRA, the total memory cost is substantially larger. RoSA, despite employing techniques for efficient computation, still requires far more memory than LoRA. Notably, we also use RoSA to demonstrate that quantizing trainable parameters (Gholami et al., 2022; Dettmers et al., 2022, 2024) with mixed-precision training (Micikevicius et al., 2018) saves significantly more memory than simply reducing trainable parameters, as shown by the memory difference between RoSA and RoSA-bf16 in Table 1 (see also related works on quantization (Dettmers et al., 2022; Guo et al., 2024a; Li et al., 2024; Dettmers et al., 2024)).
These results highlight that the primary driver of memory usage in PEFT methods is often the complexity of their computation graph and intermediate caching, rather than the number of trainable parameters alone. LoRA and its variants often rely on dropout to mitigate overfitting, which significantly contributes to memory consumption. On the other hand, SFT methods, which typically do not rely on dropout, face challenges in efficiently computing unstructured matrices, further adding to their memory demands.
@@ 4 Our Method
To address the challenges mentioned above, we propose Structured-Pruning-based Sparse Fine-Tuning (SPruFT), as illustrated in Figure 1. This is a novel approach designed to streamline computation graphs and eliminate the need for dropout layers. This method ensures memory efficiency while maintaining competitive fine-tuning performance.
Figure 1: The illustration of SPruFT: we evaluate the importance score for each neuron to select the fine-tuning indices. Then we construct the lower-dimensional fine-tuning parameter matrix ${\mathbf{W}}_{f}$.
@@ 4.1 Proposed Method
Our method selects the top-$r$ important neurons based on an importance score ${\bm{\eta}}$, where $r$ is determined by the desired number of fine-tuning parameters. It follows that the choice of importance matric becomes crucial, which we discuss in Section 4.2. Let the top $r$ neuron indices be $i_{1},i_{2},\dots,i_{r}$. After obtaining ${\bm{\eta}}$, we next construct a lower-dimensional parameter matrix ${\mathbf{W}}_{f}\in\mathbb{R}^{r\times d_{in}}$, with the row selection matrix ${\mathbf{M}}_{i_{j}j}=1$ for all $j\in[r]$ and zeros elsewhere. Using notations from Section 2, we initialize ${\mathbf{W}}_{f}$ to zero and define the final parameters $\hat{{\mathbf{W}}}$ as Equation 1 where ${\mathbf{W}}_{s}={\mathbf{M}}{\mathbf{W}}_{f}$.
@@
Let us now examine how to implement the forward to make backpropagation memory-efficient [^2] While updating the corresponding rows of ${\mathbf{W}}$ is the most efficient training, updating ${\mathbf{W}}_{s}$ provides more flexibility for adapting multiple tasks, see discussion in LoRA (Hu et al., 2022).. If the computation graph were to pass through ${\mathbf{W}}+{\mathbf{M}}{\mathbf{W}}_{f}$ (as a naïve implementation would), the gradients would be computed for all $d_{in}\times d_{out}$ parameters, which is redundant. Instead, we use the additivity of the forward function: we have, analogous to the setting of LoRA,
$f(\hat{{\mathbf{W}}},{\mathbf{x}})$ $=f({\mathbf{W}}+{\mathbf{M}}{\mathbf{W}}_{f},{\mathbf{x}})=f({\mathbf{W}},{\mathbf{x}})+f({\mathbf{M}}{\mathbf{W}}_{f},{\mathbf{x}}),$ (5)
@@
As ${\mathbf{W}}$ remains frozen during fine-tuning, backpropagation only needs to keep track of the derivatives of the second term on the RHS. In addition, ${\mathbf{M}}$ is now a fixed matrix, so the only trainable parameters are those in ${\mathbf{W}}_{f}$ and $f({\mathbf{W}}_{f},{\mathbf{x}})$ will not be cached, while LoRA requires the cache of $f({\mathbf{A}},{\mathbf{x}})$ for computing $\frac{\partial{\mathbf{h}}}{\partial{\mathbf{B}}}$ (backpropagation, Rumelhart et al. 1986). Besides, as an SFT framework, our method does not rely on any dropout layer, which also saves a huge amount of memory. We explain this in detail in Appendix C.5 and we show that the benefits in terms of memory cost are significant in Section 6.1.
An important strength of our approach is its flexibility: it can easily incorporate any desired choice of importance metrics. On the other end, it can also incorporate new ideas in PEFT research. For example, quantization (QLoRA, Dettmers et al. 2024), parameter sharing (VeRA, Kopiczko et al. 2024), and combining SFT with LoRA (RoSA, Nikdan et al. 2024) can be used.
4.2 Importance Metrics

### v2

# An Efficient Sparse Fine-Tuning with Low Quantization Error via Neural Network Pruning

@@ Abstract
Fine-tuning is an important step in adapting foundation models such as large language models to downstream tasks. To make this step more accessible to users with limited computational budgets, it is crucial to develop fine-tuning methods that are memory and computationally efficient. Sparse Fine-tuning (SpFT) and Low-rank adaptation (LoRA) are two frameworks that have emerged for addressing this problem and have been adopted widely in practice. In this work, we develop a new SpFT framework, based on ideas from neural network pruning. At a high level, we first identify "important" neurons/nodes using feature importance metrics from network pruning (specifically, we use the structural pruning method), and then perform fine-tuning by restricting to weights involving these neurons. Experiments on common language tasks show our method improves SpFT’s memory efficiency by 20–50% while matching the accuracy of state-of-the-art methods like LoRA’s variants.
@@ 1 Introduction
Despite the success, fine-tuning a model with billions of parameters requires access to heavy computational resources, even when the task datasets are fairly small. Fortunately, studies (e.g., Panigrahi et al. (2023) and references therein) show that fine-tuning only a small fraction of parameters can be effective. Parameter-efficient fine-tuning (PEFT) methods have thus been proposed to carry out this idea and address the challenge of making fine-tuning more accessible (Lialin et al., 2023). A leading PEFT approach, Low-Rank Adaptation (LoRA, Hu et al. 2022), achieves memory efficiency by simply making low-rank updates to the weight matrices in the different layers. Another class of PEFT methods is sparse fine-tuning (SpFT, Sung et al. 2021; Guo et al. 2021; Ansell et al. 2022; Nikdan et al. 2024), which learns a sparse matrix, typically an unstructured one, for updating the pre-trained weights. However, SpFT typically incurs higher memory costs than LoRA during the fine-tuning process, because of the unstructured sparsity. Several works aim to mitigate the memory complexity of SpFT (Mofrad et al., 2019; Holmes et al., 2021; Nikdan et al., 2023; 2024), often at the cost of increased running time and more complex implementations of sparse kernels. Besides PEFTs, techniques like Zeroth-Order optimization (Malladi et al., 2023; Guo et al., 2024b) and quantization (Gholami et al., 2022; Dettmers et al., 2022; 2024) can further enhance memory and training efficiency for fine-tuning, including LoRA and SpFT.
As LLMs increase in scale, advancing efficient sparse matrix computation, PEFT, and efficient training remains a crucial problem. Towards this goal, we study the question: Can sparse fine-tuning be improved to create a memory- and parameter-efficient framework, while avoiding additional implementations of sparse operations and without increasing the training time complexity? We answer this question in the affirmative, by proposing a new SpFT framework for fine-tuning LLMs that achieves memory- and parameter-efficiency while maintaining or even improving performance on downstream tasks. Our approach utilizes NN pruning techniques to identify a subset of fine-tuning parameters and employs a matrix decomposition-based computation for efficient fine-tuning. This design enables the integration of ideas from model compression, SpFT, and matrix decomposition methods.
@@ 1.1 Our Contributions
  • We leverage ideas from network pruning to improve SpFT, achieving a significant memory efficiency considerably lower than that of the popular LoRA. Our method uses only standard tensor operations, eliminating the need for custom sparse tensor operations. Additionally, our approach supports fine-tuning quantized base models to further reduce memory footprints.
  • We analyze the memory assignment of several PEFT methods and suggest that computation graphs can affect memory more significantly than the number of trainable parameters. In addition, we validate our methods in various fine-tuning tasks and provide practical guidance on training strategies.
  • We propose two variants of the Taylor importance for different settings in image and language tasks: class-aware Taylor and Zeroth-Order Taylor. The first is designed for tasks where class-wise accuracy is important (in addition to overall accuracy), such as image classification. Zeroth-Order Taylor is designed for large language models and requires memory only equal to that of a forward pass. In addition, we show how to effectively reduce the estimation variance of the Zeroth-Order estimator.
@@ 2 Background and Related Work
Parameter-Efficient and Memory-Efficient Fine-Tuning: In various language and vision tasks, the “pre-train then fine-tune” paradigm has been shown highly effective. PEFT methods (Lialin et al., 2023) fine-tune a small subset of the parameters of a large pre-trained model in order to accelerate the training process. We begin by introducing SpFT and LoRA, two popular approaches for PEFT.
Sparse Fine-Tuning: SpFT formulates the fine-tuning process as learning another weight matrix ${\bm{\delta}}$:
$\hat{{\mathbf{W}}}={\mathbf{W}}+{\bm{\delta}},$ (1)
${\mathbf{h}}=f(\hat{{\mathbf{W}}},{\mathbf{x}})=f({\mathbf{W}}+{\bm{\delta}},{\mathbf{x}}),$ (2)
where ${\mathbf{h}}\in\mathbb{R}^{d_{out}}$ and ${\mathbf{x}}\in\mathbb{R}^{d_{in}}$ are the input and output of the hidden layer, respectively, $f(\cdot)$ is the forward function, ${\mathbf{W}}\in\mathbb{R}^{d_{out}\times d_{in}}$ represents the frozen pre-trained parameters, and $\hat{{\mathbf{W}}}\in\mathbb{R}^{d_{out}\times d_{in}}$ denotes the final parameters used during inference for the fine-tuning task. The matrix ${\bm{\delta}}\in\mathbb{R}^{d_{out}\times d_{in}}$ is sparse and is initialized at ${\bm{0}}$. Such a decomposition is done for every layer in the neural network. SpFT methods try to limit the number of parameters to fine-tune. For selecting non-zero indices, Diff pruning (Guo et al., 2021) learns a mask for ${\bm{\delta}}$ (using a standard Backprop algorithm), while FISH Mask (Sung et al., 2021) uses Fisher information to identify important indices in ${\mathbf{W}}$. Lottery Ticket SpFT (Ansell et al., 2022) fine-tune the whole model for one epoch, then use ${\bm{\delta}}$ itself as an importance score to decide which parameters to fine-tune subsequently. Robust Adaptor (RoSA, Nikdan et al. 2024) combines the above SpFTs with LoRA and outperforms all these approaches. However, the key challenge of all SpFT methods is that they do not sufficiently reduce memory usage, as ${\bm{\delta}}$ keeps the dimensionality of ${\mathbf{W}}$, and thus standard libraries do not yield memory improvements.
@@
However, these techniques come with trade-offs, particularly in terms of time complexity and implementation complexity. Achieving memory efficiency often requires a significant increase in time complexity. To mitigate this, some approaches employ optimizations implemented in C++ or lower-level languages, such as those used in (Gale et al., 2020; Nikdan et al., 2023; 2024), to accelerate the training process.
Low-Rank Adaptation (LoRA): Instead of requiring ${\bm{\delta}}$ to be sparse, low-rank adaptation aims to find update matrices that are of small rank:
@@
where $\alpha$ is the LoRA scaling hyper-parameter, ${\mathbf{B}}\in\mathbb{R}^{d_{out}\times r},\ {\mathbf{A}}\in\mathbb{R}^{r\times d_{in}}$ are the low-rank matrices with $r\ll d_{in},d_{out}$. During inference, the ${\mathbf{B}}{\mathbf{A}}$ term can be merged into ${\mathbf{W}}$ to maintain the inference latency of the original model. During training, owing to the fact that $f$ is additive for both the self-attention blocks and the subsequent feed-forwarding neworks (FFN) of transformers (Vaswani, 2017), backpropagation can be performed efficiently for the ${\mathbf{B}},{\mathbf{A}}$ parameters. Due to LoRA’s simplicity and effectiveness, numerous variants have been proposed to enhance the performance, e.g., QLoRA (Dettmers et al., 2022; Guo et al., 2024a; Li et al., 2024; Dettmers et al., 2024), DoRA (Liu et al., 2024), RoSA (Nikdan et al., 2024), and VeRA (Kopiczko et al., 2024). These methods have achieved exceptional performance, often comparable to full fine-tuning across a range of tasks.
@@ 3 Number of Trainable Parameters Is Not Everything
Before introducing our approach, we want to emphasize that in PEFT research, reducing the number of trainable parameters is not the most critical factor for minimizing memory consumption. While certain PEFT methods explicitly aim to lower the number of trainable parameters to reduce memory usage, the impact of this reduction diminishes once the parameter count is sufficiently small. To investigate this further, we compare several representative methods of SpFT and low-rank methods, focusing on their memory footprints during training, as shown in Figure 1. For precisely, we assume full-precision training to exclude memory costs introduced by quantization or other compression techniques and implementations. Notably, quantizing the model to 4-bit precision can yield an additional memory savings of approximately 75%.
During neural network training, backpropagation requires caching a large number of intermediate activations to compute gradients efficiently. The memory cost of these intermediate values is largely influenced by the structure of the computation graph. When the number of trainable parameters is small, the memory consumed by intermediate activations (see green bars in Figure 1) often dominates memory usage apart from the model weights.
Among the LoRA-based methods, VeRA attempts to reduce memory by sharing a pair of low-rank matrices across layers, thereby reducing the number of trainable parameters. However, as shown in Figure 1, this results in only marginal memory savings—around 0.5GB to 2GB depending on the maximum token length, which is almost negligible. In contrast, DoRA and RoSA incur significantly higher memory usage due to their more complex computation graphs and reliance on unstructured sparse matrices. For instance, DoRA decomposes LoRA’s matrices into separate magnitude and direction components (see Figure 4 in Appendix D.4), which substantially increases memory requirements for activation caching. While DoRA’s trainable parameter cost is similar to that of LoRA, its overall memory consumption is considerably higher. RoSA also consumes much more memory than LoRA despite incorporating efficiency-oriented design choices. These findings suggest that a simple computation graph can be a far more significant contributor to memory usage than reducing trainable parameters.
Figure 1: Memory footprints (in GB) of fine-tuning full precision Llama2 using LoRA (Hu et al., 2022), RoSA (Nikdan et al., 2024), DoRA (Liu et al., 2024), VeRA (Kopiczko et al., 2024), and ours. We set $r=32,d=1.2\%$ for RoSA, $r=128$ for ours, and $r=64$ for the others. Note that RoSA has its own official implementation, which may influence memory consumption, whereas LoRA, DoRA, and VeRA are integrated into the PEFT library provided by Hugging Face. More details please see Appendix D.3.
@@ 4 Our Method
To address the challenges mentioned above, we propose Structured-Pruning-based Sparse Fine-Tuning (SPruFT), as illustrated in Figure 2. This is a novel approach designed to streamline computation graphs and eliminate the need for implementing sparse tensor operations. This method ensures memory efficiency while maintaining competitive fine-tuning performance.
Figure 2: The illustration of SPruFT: we evaluate the importance score for each neuron to select the fine-tuning indices. Then we construct the lower-dimensional fine-tuning parameter matrix $\mathrm{\Delta}{\mathbf{W}}$.
@@ 4.1 Proposed Method
Our method selects the top-$r$ important neurons based on an importance score ${\bm{\eta}}$, where $r$ is determined by the desired number of fine-tuning parameters. It follows that the choice of importance matric becomes crucial, which we discuss in Section 4.3. Let the top $r$ neuron indices be $i_{1},i_{2},\dots,i_{r}$. After obtaining ${\bm{\eta}}$, we next construct a lower-dimensional parameter matrix $\mathrm{\Delta}{\mathbf{W}}\in\mathbb{R}^{r\times d_{in}}$, with the row selection matrix ${\mathbf{M}}_{i_{j}j}=1$ for all $j\in[r]$ and zeros elsewhere. Using notations from Section 2, we initialize $\mathrm{\Delta}{\mathbf{W}}$ to zero and define the final parameters $\hat{{\mathbf{W}}}$ as Equation 1 where ${\bm{\delta}}={\mathbf{M}}\mathrm{\Delta}{\mathbf{W}}$.
@@
Let us now examine how to implement the forward to make backpropagation memory-efficient [^2] While updating the corresponding rows of ${\mathbf{W}}$ is the most efficient training, updating ${\bm{\delta}}$ provides more flexibility for adapting multiple tasks, see discussion in LoRA (Hu et al., 2022).. If the computation graph were to pass through ${\mathbf{W}}+{\mathbf{M}}\mathrm{\Delta}{\mathbf{W}}$ (as a naïve implementation would), the gradients would be computed for all $d_{in}\times d_{out}$ parameters, which is redundant. Instead, we use the additivity of the forward function: we have, analogous to the setting of LoRA,
$f(\hat{{\mathbf{W}}},{\mathbf{x}})$ $=f({\mathbf{W}}+{\mathbf{M}}\mathrm{\Delta}{\mathbf{W}},{\mathbf{x}})=f({\mathbf{W}},{\mathbf{x}})+f({\mathbf{M}}\mathrm{\Delta}{\mathbf{W}},{\mathbf{x}}),$ (5)
@@
As ${\mathbf{W}}$ remains frozen during fine-tuning, backpropagation only needs to keep track of the derivatives of the second term on the RHS. In addition, ${\mathbf{M}}$ is now a fixed matrix, so the only trainable parameters are those in $\mathrm{\Delta}{\mathbf{W}}$ and $f(\mathrm{\Delta}{\mathbf{W}},{\mathbf{x}})$ will not be cached, while LoRA requires the cache of $f({\mathbf{A}},{\mathbf{x}})$ for computing $\frac{\partial{\mathbf{h}}}{\partial{\mathbf{B}}}$ (backpropagation, Rumelhart et al. 1986). Besides, as an SpFT framework, our method does not rely on any dropout layer, which also saves a huge amount of memory. We explain this in detail in Appendix D.5 and we show that the benefits in terms of memory cost are significant in Section 6.1.
4.2 QSPruFT: Extension Approach with Low Quantization Error
Our approach is model-agnostic, allowing users to integrate recent advances in PEFT such as QLoRA (Dettmers et al., 2024), LoftQ (Li et al., 2024), and QPiSSA (Meng et al., 2024). QLoRA quantizes the base model to Normal Float 4-bit (NF4) and fine-tunes full-precision LoRA matrices ${\mathbf{A}}$ and ${\mathbf{B}}$. Based on QLoRA, LoftQ and QPiSSA propose alternative initialization strategies for ${\mathbf{A}}$ and ${\mathbf{B}}$ to reduce quantization error. This error is defined as:
${\mathbf{W}}^{res}={\mathbf{W}}-{\mathbf{W}}^{\text{NF4}},$ (6)
and both LoftQ and QPiSSA use singular value decomposition (SVD) to initialize ${\mathbf{A}}$ and ${\mathbf{B}}$ such that ${\mathbf{B}}^{init}{\mathbf{A}}^{init}\approx{\mathbf{W}}^{res}$.
In our method, when applied to quantized base models, we do not require decomposition to approximate ${\mathbf{W}}^{res}$. Since we fine-tune only selected rows, we can directly initialize $\mathrm{\Delta}{\mathbf{W}}$ using the corresponding rows from ${\mathbf{W}}^{res}$, where the quantization error of the fine-tuning rows is zero. This offers a potential advantage over QLoRA, QDoRA, LoftQ, and QPiSSA in terms of accuracy.
4.3 Broader Impacts: Importance Metrics